OpenStudy (anonymous):

tell you what operation to use.6. Once there was Honda XR 100. It was for sale. It was $1,705. My Mom said, "Maybe son, I will think about it." I said, "Mom, I only have $400." How much more money do I need?

OpenStudy (anonymous):

Subtraction..

OpenStudy (anonymous):

ok the price of the vehicle is 1705 you have 400 how much more money do you need (call it x) 400 + x = 1705 x = 1705 - 400 x = $1305

OpenStudy (anonymous):

were key word

OpenStudy (anonymous):

?

OpenStudy (anonymous):

for Subtraction.